# PFDial: A Structured Dialogue Instruction Fine-tuning Method Based on UML Flowcharts

### v1

@@ Abstract
Ming Zhang1thanks:   Equal Contribution.   Yuhui Wang1∗,   Yujiong Shen1,   Tingyi Yang1,   Changhao Jiang1,   Yilong Wu1, Shihan Dou1,   Qinhao Chen1,4,   Zhiheng Xi1,   Zhihao Zhang1,   Yi Dong1,   Zhen Wang2,   Zhihui Fei2, Mingyang Wan2,   Tao Liang2,   Guojun Ma2thanks:   Corresponding Author.,   Qi Zhang1†,   Tao Gui1,3,   Xuanjing Huang1 1 School of Computer Science, Fudan University 2 Douyin Co., Ltd. 3 Institute of Modern Languages and Linguistics, Fudan University 4 Graduate School of Arts and Sciences, Columbia University {mingzhang23,wangyuhui21}@m.fudan.edu.cn maguojun@bytedance.com qz@fudan.edu.cn
@@ 2.2 Dataset Construction Process
2.2.1 Flow Chart Collection
@@
2.2.2 Textual Representation Conversion
@@
2.2.3 State Transition Information Extraction
@@ 3.2 Main Results
Table 2 presents our detailed experimental results on the main experiments. Our experiments demonstrate significant improvements over baseline models across all parameter scales. Even our smallest 0.5B parameter model achieves 91.35% accuracy on out-of-domain tests, with particularly strong performance on decision branches. Our 8B model achieves sota performance with 97.29% accuracy on out-of-domain tests, surpassing GPT-4o by 11%. In decision branches, our 8B parameter model achieves 96.88% accuracy, surpassing GPT-4o by 43.88%.
Figure 4: Results on Model Performance with Different Formats: The left plot shows the sequential accuracy for different models across three different formats (Format A, Format B, and Format C). The right plot shows the decision accuracy for the same models under the same formats.
@@ 3.5 Format Abliation Studies
We conducted experiments for three state representation formats: Format A: natural language description (default method), Format B: state codes (e.g., S1, C2), and Format C: combining codes and descriptions to explore the impact of different data formats on model performance and the underlying reasons. Specific cases of data in different formats and the visualization results can be seen in Appendix C.3. To ensure fairness, we fine-tuned the base model on data in all three formats and tested it with corresponding test sets containing the same content.
@@
In particular, $head\_layer_{27}\_head_{20}$ is crucial in all three formats. We visualized this attention head’s scores for all three formats in Figure 6, Figure 7, and Figure 8, respectively, focusing on significant differences. The attention scores for both the natural language and hybrid formats were concentrated at the intersection of the user’s current state and the corresponding state in the flowchart. Format C, which includes some state codes, showed a more moderated concentration. In contrast, Format B did not exhibit such clustering. We hypothesize that the introduction of state codes allows the model’s attention to generalize across different input parts, rather than being confined to specific segments. This enables models to better understand user inputs and facilitates learning of global logic, such as condition checking and state selection. This also reasonably explains previous results.
@@ 5 Conclusion
Our research presents a comprehensive set of experiments, yet it is not without limitations. First, the Chinese-centric nature of our dataset introduces potential cross-lingual generalization constraints.Second, the scarcity of standardized flowchart benchmarks in Chinese process specifications increases the risk of domain-specific biases, despite our rigorous validation framework; while the potential residual inconsistencies in flow-to-text conversion may emerge from the inherent subjectivity in interpreting semantic structures.
@@ C.2 Supplementary Data for Data Scaling Experiment
While the section 3.3 presents the overall accuracy trend with increasing training samples, here we provide the complete experimental results. Tables 11 and 12 present the performance results for Strategy A and Strategy B respectively. Each table shows overall accuracy, decision accuracy, and sequential accuracy metrics on both ID and OOD tests across different training sample sizes. The information of dataset with different training sample sizes is shown in Table 10.
@@ C.3 Details in Format Ablation Study
In this section, we present specific cases of data in different formats and the visualization results for $head\_layer_{27}\_head_{20}$. As shown in Figure 6, 7, and 8, we visualize the attention patterns for Format A (Natural Language) in Table 13, Format B (State Code) in Table 14, and Format C (Hybird) in Table 15 respectively.
@@ 3. Ensure that the syntax of PlantUML is correct
(a) Format A
Figure 6: local attention score of $head\_layer_{27}\_head_{20}$ using Format A
(b) Format B
Figure 7: local attention score of $head\_layer_{27}\_head_{20}$ using Format B
(c) Format C
Figure 8: local attention score of $head\_layer_{27}\_head_{20}$ using Format C
@@
Table 13: Example Data Display - Format A
@@
Table 14: Example Data Display - Format B
@@
Table 15: Example Data Display - Format C

### v2

@@ Abstract
Ming Zhang1thanks:   Equal Contribution.,   Yuhui Wang1∗,   Yujiong Shen1∗,   Tingyi Yang1,   Changhao Jiang1, Yilong Wu1,  Shihan Dou1,   Qinhao Chen5,   Zhiheng Xi1,   Zhihao Zhang1,   Yi Dong1, Zhen Wang2,   Zhihui Fei2,   Mingyang Wan2,   Tao Liang2, Guojun Ma2thanks:   Corresponding Author.,   Qi Zhang1,4†,   Tao Gui3,4,   Xuanjing Huang1,3,4 1 Computation and Arificial intelligence innovative College, Fudan University 2 Douyin Co., Ltd. 3 Institute of Modern Languages and Linguistics, Fudan University 4 Institute of Trustworthy Embodied Artificial Intelligence, Fudan University 5 Graduate School of Arts and Sciences, Columbia University mingzhang23@m.fudan.edu.cn maguojun@bytedance.com qz@fudan.edu.cn
@@ 2.2 Dataset Construction Process
Flow Chart Collection
@@
Textual Representation Conversion
@@
State Transition Information Extraction
@@ 3.2 Main Results
Table 2 presents our detailed experimental results on the main experiments. Our experiments demonstrate significant improvements over baseline models across all parameter scales. Even our smallest 0.5B parameter model achieves 91.35% accuracy on out-of-domain tests, with particularly strong performance on decision branches. Our 8B model achieves sota performance with 97.29% accuracy on out-of-domain tests, surpassing GPT-4o by 11%. In decision branches, our 8B parameter model achieves 96.88% accuracy, surpassing GPT-4o by 43.88%. See Appendix E for a detailed case study analyzing the reasons behind model errors.
Figure 4: Results on Model Performance with Different Formats: The left plot shows the sequential accuracy for different models across three different formats (Format-NL, Format-SC, and Format-Hybrid). The right plot shows the decision accuracy for the same models under the same formats.
@@ 3.5 Format Abliation Studies
We conducted experiments for three state representation formats: Format-NL: natural language description (default method), Format-SC: state codes (e.g., S1, C2), and Format-Hybrid: combining codes and descriptions to explore the impact of different data formats on model performance and the underlying reasons. Specific cases of data in different formats and the visualization results can be seen in Appendix C.3. To ensure fairness, we fine-tuned the base model on data in all three formats and tested it with corresponding test sets containing the same content.
@@
In particular, $head\_layer_{27}\_head_{20}$ is crucial in all three formats. We visualized this attention head’s scores for all three formats in Figure 6, Figure 7, and Figure 8, respectively, focusing on significant differences. The attention scores for both the natural language and hybrid formats were concentrated at the intersection of the user’s current state and the corresponding state in the flowchart. Format-Hybrid, which includes some state codes, showed a more moderated concentration. In contrast, Format-SC did not exhibit such clustering. We hypothesize that the introduction of state codes allows the model’s attention to generalize across different input parts, rather than being confined to specific segments. This enables models to better understand user inputs and facilitates learning of global logic, such as condition checking and state selection. This also reasonably explains previous results.
@@ 5 Conclusion
Our research presents a comprehensive set of experiments, yet it is not without limitations. First, the Chinese-centric nature of our dataset introduces potential cross-lingual generalization constraints. Second, the scarcity of standardized flowchart benchmarks in Chinese process specifications increases the risk of domain-specific biases, despite our rigorous validation framework. Besides, the potential residual inconsistencies in flow-to-text conversion may emerge from the inherent subjectivity in interpreting semantic structures.
Acknowledgments
The authors wish to thank anonymous reviewers for their helpful comments. This work was partially funded by National Natural Science Foundation of China (No.62376061,62206057), Shanghai Rising-Star Program (23QA1400200), Natural Science Foundation of Shanghai (23ZR1403500), Program of Shanghai Academic Research Leader under grant 22XD1401100.
@@ C.2 Supplementary Data for Data Scaling Experiment
While the section 3.3 presents the overall accuracy trend with increasing training samples, here we provide the complete experimental results. Tables 13 and 14 present the performance results for Strategy A and Strategy B respectively. Each table shows overall accuracy, decision accuracy, and sequential accuracy metrics on both ID and OOD tests across different training sample sizes. The information of dataset with different training sample sizes is shown in Table 12.
@@ C.3 Details in Format Ablation Study
In this section, we present specific cases of data in different formats and the visualization results for $head\_layer_{27}\_head_{20}$. As shown in Figure 6, 7, and 8, we visualize the attention patterns for Format-NL (Natural Language) in Table 15, Format-SC (State Code) in Table 16, and Format-Hybird in Table 17 respectively.
@@ 3. Ensure that the syntax of PlantUML is correct
Appendix E Case Study
We analyzed specific cases to provide intuitive explanations in Table 10 and 11. First, generally speaking, these models fail because they don’t precisely provide complete descriptions of the next state, or they generate states related to user input but not in the original flowchart, or they don’t follow the format required in the system prompt. Our task examines models’ ability to predict the next state in sequential branches and decision branches. Below we provide examples and analyze why GPT-4o performs poorly in these areas.
PlantUML @startuml start :Customer arrives at photo shop; :Submit photo files for printing; :Select print size and quantity; :Choose paper quality and surface effect; :Confirm print order and price; if (Photo quality check passed?) then (yes)         :Pay fee;         :Photo printing in progress;         if (Printing completed?) then (yes)               :Display printed photos;               :Customer satisfied?;               if (Customer satisfied?) then (yes)                     :Complete transaction;               else (no)                     :Reselect photos;                     :Reprint;                     :Complete transaction;               endif         else (no)               :Wait for photo printing;               :Repeatedly check if printing is complete;         endif else (no)         :Notify poor photo quality;         :Reselect photos;         :Reprint;         :Complete transaction; endif :Customer leaves the photo shop; stop @enduml
Current State Repeatedly check if printing is complete
User Input Printing is completed.
Predicted Next State Display printed photos (Incorrect)
Correct Next State Customer leaves the photo shop (Correct)
Table 10: Case Study in Sequential Branch
E.1 Sequential Branch
For sequential branches, the model only need to capture the current state and state transition. This is typically simple because adjacent states in PlantUML are usually adjacent lines. However, when multiple decision branches are nested, things become complicated and the line distance between adjacent states increases. For example, in Table 10, after "Repeatedly check if printing is complete", all nested loops have ended and should transition to "Customer leaves the photo shop". However, due to the high complexity caused by multiple nested sequential branches, GPT-4o struggles to accurately capture the contextual state relationships, resulting in errors.
E.2 Decision Branch
For decision branches, the model needs to precisely transition to the next state based on the current state and user input. Unfortunately, in most wrong cases, due to the high similarity between user input and the first state of the correct branch during condition judgment, the model tends to skip this first state and directly enter subsequent states, failing to accurately follow the agreed path. More challenging issues arise when decision branch conditions involve more than just yes/no decisions, requiring models to comprehensively understand user input and judge their true intentions. For example, in Table 11, when the user inputs "The position of the lighting fixture needs to be adjusted.", it indicates that the lighting position needs to be modified at this point. However, the model struggles to strictly adhere to the process constraints and makes an incorrect state prediction based on the user input, skipping the correct state of this branch.
PlantUML @startuml start :Communicate lighting design plan with construction team; :Explain design requirements and fixture layout; :Provide technical support and guidance; :Coordinate installation time and progress; :Supervise fixture installation process; :Perform fixture debugging and brightness adjustment; repeat :Confirm fixture layout and installation position; :Supervise fixture installation process; repeat while(Installation and debugging unsuccessful) :Check fixture installation quality and safety; if (Need to adjust fixture position?) then (yes)         :Negotiate adjustment plan;         :Reinstall or adjust fixture position; else (no)         :Confirm final installation result; endif :Final acceptance of lighting system; :Ensure lighting system meets design requirements; :Complete construction documents and records; stop @enduml
Current State Need to adjust fixture position?
User Input The position of the lighting fixture needs to be adjusted.
Predicted Next State Reinstall or adjust fixture position (Incorrect)
Correct Next State Negotiate adjustment plan (Correct)
Table 11: Case Study in Decision Branch
@@ E.2 Decision Branch
(a) Format-NL
Figure 6: local attention score of $head\_layer_{27}\_head_{20}$ using Format-NL
(b) Format-SC
Figure 7: local attention score of $head\_layer_{27}\_head_{20}$ using Format-SC
(c) Format-Hybrid
Figure 8: local attention score of $head\_layer_{27}\_head_{20}$ using Format-Hybrid
@@
Table 15: Example Data Display - Format-NL
@@
Table 16: Example Data Display - Format-SC
@@
Table 17: Example Data Display - Format-Hybrid